OpenStudy (anonymous):

(-3b + 5x) + (-2b + 3x + 5)

OpenStudy (mathstudent55):

You are adding the polynomials, so the parentheses are unnecessary. Rewrite the expression with no parentheses.

OpenStudy (anonymous):

Combine like terms so -3b + (-2b)= -5b 5x + 2x= 7x and since '5' has no like term in the other equation it just stays 5 so put the terms together and the final answer should be: -5b + 7x + 5

OpenStudy (mathstudent55):

Then the next step is to combine like terms.

OpenStudy (anonymous):

you think you can help me on a quiz

OpenStudy (mathstudent55):

@ChihuahuaAdorer Great job above, but notice the x term in the second polynomial is 3x not 2x.

OpenStudy (anonymous):

can u plz help me after??? @mathstudent55

OpenStudy (anonymous):

ah yes sorry so the final answer is actually -5b + 8x +5

OpenStudy (anonymous):

(9x + 6b) + (-2x - 2b - 1)

OpenStudy (anonymous):

Okay so I will help you along on this one but I want you to try to do this. So add like terms.

OpenStudy (anonymous):

-11+8b-1

OpenStudy (anonymous):

Very close! But there is a negative in front of the the 2x and the 2b so: 9x + (-2x) would be 7x because you have to subtract because the negative is there. and 6b + (-2b) would equal 4b because you also have to subtract because the negative in front of the two. The final equation would be 7x + 4b - 1

OpenStudy (anonymous):

4x3 - 3x) + (-5x3 + 4x - 5)

OpenStudy (anonymous):

is it \[4x ^{3} \] and \[-5x ^{3}\] ?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

okay so combine like terms again

OpenStudy (anonymous):

-9 6-5

OpenStudy (anonymous):

Remember that the negatives on the equation to the left cause you to subtract not add! :)

OpenStudy (anonymous):

-9+6x-5

OpenStudy (anonymous):

\[(4x ^{3}-3x) + (-5x ^{3} + 4x -5)\] \[(4x ^{3} + (-5x ^{3})) + (-3x + 4x) -5\] ^this is where you are combining like terms \[-1x ^{3} + 1x -5 (or) x ^{3} + x -5\] ^and after combining the terms your equation should look like this

OpenStudy (anonymous):

\[-x ^{3} + x -5\] is the final answer

OpenStudy (anonymous):

7n3 - 9n + 7) + (-8n3 + 6n)

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

idk im going to skip it i want to do the test and its like ten questions

OpenStudy (anonymous):

do you understand how to do these type of equations though? if you need anymore help just tag me! :)

OpenStudy (anonymous):

@ChihuahuaAdorer (-4ab+8a-2) - (-4ab-9)

OpenStudy (anonymous):

Combine like terms: (-4ab- (-4ab))+(8a-0)+ (-2- (-9)) (-4ab- (-4ab))= 0 (8a-0)= 8a (-2-(-9))=7 So 8a+7 should be your final answer